Q-Find the equation of circle passing through the origin, radius 17 and ordinate of the centre is -15.Please show the solution

Grade:11

1 Answers

Har Simrat Singh
11 years ago

Let centre be h,k given k= -15 also radius is 17 and passes thru origin therefore distance of centre to origin is radius ie

17= sqrt( h2 + (-15)2) or 289 = h2 + 225

from here h= +-8 

hence equation is (x-8)2 + (y+15)2 = 289 or (x+8)2+(y+15)2=289